Whats the equation of a line with a slope of 2/3 and a y-intercept of (0, -2)?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

Whats the equation of a line with a slope of 2/3 and a y-intercept of (0, -2)?
--------------
Ans: y = (2/3)x -2
---------------------
Cheers,
Stan H.
--------------

Answer by richwmiller(17219)

y=mx+b
where m is the slope
b is the y intercept when x=0
y=2/3*x-2